tom is four times as old as Todd in five years Tom's age will be 1 year less than three times as old as Todd. how old is Tom ?
Answer by lwsshak3(11628)

tom is four times as old as Todd in five years Tom's age will be 1 year less than three times as old as Todd. how old is Tom ?
***
let x=todd's age
4x=tom's age
..
In 5 yrs:
x+5=todd.s age
4x+5=tom's age
..
4x+5=3(x+5)-1(1 yr less than 3 times todd's age)
4x+5=3x+15-1
x=9
4x=36
how old is Tom ? 36